Q2. MNO Inc., a publicly traded manufacturing firm in the United States, has provided the following financial information in its application for a loan. All numbers are in thousands of dollars Assets Cash Accounts receivables Inventory Liabilities and Equity Accounts payable Notes payable Accruals Long-term debt Equity (ret. earnings $22 400 Total liabilities and equityS700 S 20 90 90 S 30 90 30 Plant and equipment Total assets 500 S700 Also assume sales $500,000 cost of goods sold $360,000; and the market value of equity is equal to the book value a. What is the Altman discriminant function value for MNO Inc.? Recall that Net working capital Current assets Current liabilities Current assets = Cash + Accounts receivable + inventories. Current liabilities- Accounts payable + Accruals + Notes payable EBIT-Revenues Cost of goods sold Altmans discriminant function is given by: Z-1.2X1+ 1.4X2 + 3.3X3 + 0.6X4+ /.0X All numbers are in $000s b.Based on the Altmans Z-score only, should you approve MNO Inc.s application to your bank for a S500,000 capital expansion loan?
Answer #1
a) Altman's Z score is given by the formula:
Z-Score = ([Working Capital / Total Assets] x 1.2) + ([Retained Earnings / Total Assets] x 1.4) + ([Operating Earnings / Total Assets] x 3.3) + ([Market Capitalization / Total Liabilities] x 0.6) + ([Sales/ Total Assets] x 1.0)
Workings:
Working capital = 200000-150000 = 50000 Ratio Weightage Weighted Score
Working capital/Total assets = 50000/700000 = 0.071 1.20 0.0857
Retained earnings/Total assets = 22000/700000 = 0.031 1.40 0.0440
Operating earnings/Total assets = (500000-360000)/700000 0.200 3.30 0.6600
Market capitalization/Total liabilities = 400/300 = 1.333 0.60 0.8000
Sales/Total assets = 500000/700000 = 0.714 1.00 0.7143
Z = 2.3040
b) Grading of Z Score:
If the Z score is higher than 3.0, the company is a safe company and
has no chance of bankruptcy.
If the z score is below 1.8, the probability of the company becoming
bankrupt is high.
If the score is between 1.8 and 3.0, there is a chance of becoming
bankrupt.
As the Z score of MNO Inc, is less than 3.0, there is a chance of the
company becoming bankrupt. Hence, the loan need not be given.